# Transit for All: Mapping Equitable Bike2Subway Connection using Region Representation Learning

### v1

@@ 7. Conclusion
In this study, we introduced Transit for All (TFA), a spatial computing framework designed to support equitable expansion of BSS. TFA integrates multimodal geospatial data, including demographics, points-of-interest, and satellite imagery, to accurately predict bike demand at cold-start stations, significantly outperforming traditional approaches.
Additionally, we developed a novel weighted Public Transport Accessibility Level (wPTAL) metric, combining predicted bike demand with traditional accessibility measures. Our results showed that wPTAL effectively identifies previously overlooked disparities in transit accessibility among various income and racial groups, and provides strategic recommendations that enhance transit equity.
In the future, we plan to extend our framework to include diverse transportation modes and additional urban contexts, further supporting inclusive and adaptive mobility strategies aligned with the UN Sustainable Development Goal 11.2.

### v2

@@ 7. Conclusion
In this study, we introduced Transit for All (TFA), a spatial computing framework designed to support equitable expansion of BSS. TFA integrates multimodal geospatial data, including demographics, POI, and satellite imagery, to accurately predict bike demand at cold-start stations, significantly outperforming traditional approaches. Additionally, we developed a novel weighted Accessibility Level (wAL) metric, combining predicted bike demand with traditional accessibility measures.
Our results showed that wAL effectively identifies previously overlooked disparities in transit accessibility among various income and racial groups, and provides strategic recommendations that enhance transit equity. In the future, we plan to extend our framework to include diverse transportation modes and additional urban contexts, further supporting inclusive and adaptive mobility strategies aligned with the UN Sustainable Development Goal 11.2.